OpenStudy (anonymous):

one of the solutions to the quadratic equation x2 + 3x - 28 is -7 is there another?

razor99 (razor99):

could u post the question

razor99 (razor99):

u mean u have to slove that eqaution right

OpenStudy (anonymous):

yeah factoring to be exact

OpenStudy (anonymous):

i just started learning these on my own well i want to learn so i am not sure completely whats to be donw i can give examples from my book thopugh

razor99 (razor99):

5x+is+21

razor99 (razor99):

did u get it

OpenStudy (dumbcow):

7*-4 = -28 7 -4 = 3 --> x^2 +3x -28 = (x+7)(x-4) solutions are -7 and 4